# Gromov’s Compactness Theorem for the Intrinsic Timed-Hausdorff Distance

### v2

@@ Abstract.
The intrinsic timed-Hausdorff distance between timed-metric-spaces, first introduced by Sakovich–Sormani, yields a weak notion of convergence for space-times. In this paper we prove a compactness theorem for the intrinsic timed-Hausdorff convergence of timed-metric-spaces using timed-Fréchet maps. Our proof introduces the notion of “addresses” and provides a new way of stating Gromov’s original compactness theorem for Gromov–Hausdorff convergence of metric spaces. We also obtain a new Arzelà–Ascoli theorem for real valued uniformly bounded Lipschitz functions on Gromov–Hausdorff converging compact metric spaces.
@@ 1. Introduction
One would also like to study sequences of smooth Lorentzian manifolds which are not diffeomorphic. In [26], Sakovich and Sormani introduced the notion of the intrinsic timed-Hausdorff convergence as one possible way to achieve this. They described a method of converting smooth causally null compactifiable space-times into compact timed-metric-spaces in a canonical way using the cosmological time function of Anderson–Galloway–Howard [4] and the null distance of Sormani–Vega [30]. In general, timed-metric-spaces, $(X_{j},d_{j},\tau_{j})$, are metric spaces, $(X_{j},d_{j})$, endowed with a Lipschitz one  [^1] In this manuscript we refer to Lipschitz functions with Lipschitz constant bounded above by $1$ as Lipschitz one function, $\tau_{j}\colon X_{j}\to\mathbb{R}$, and a causal structure. They then defined the intrinsic timed-Hausdorff convergence of these timed-metric-spaces.
@@
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed metric spaces that are equibounded,
@@
and have uniform bounds on their Lipschitz one functions
@@
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed metric space $(X_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance and time preserving timed-Fréchet maps,
@@
We remark that there are special classes of space-times where the cosmological time functions are easy to control. These include the big-bang space-times and future developed space-times. Sormani–Vega [29] and Sakovich–Sormani [26] defined convergence for such space-times based on pointed GH convergence of Gromov [18] and GH Convergence for Metric Pairs of Che, Galaz-García, Guijarro, and Membrillo Solis [7]. Such notions already have compactness theorems proven by Gromov [18] and by Che and Ahumada Gómez [1] respectively. The intrinsic timed-Hausdorff distance applies to a far wider class of space-times including sub-space-times that exhaust asymptotically flat space-times [26].
It should be noted that some authors have used the Sormani–Vega null distance of [30] to convert space-times into metric spaces without keeping track of the time functions. See the work of Allen, Burtscher, García-Heveling, Kunzinger, and Steinbauer in [2, 3, 6, 20]. Furthermore, Allen, Burtscher, Kunzinger, and Steinbauer [2, 3, 20] have applied Gromov’s original compactness theorem to obtain a GH limit space which is only a metric space with no causal structure unlike our timed metric spaces.
@@
Our notion of “addresses” provides a new way of stating Gromov’s Compactness Theorem for sequences of metric spaces, that we state and prove in Section 4.1. We also restate an Arzelà–Ascoli Theorem for sequences of real valued uniformly bounded Lipschitz functions on a converging sequence of metric spaces using “addresses” in Section 4.2. We suggest a more general Arzelà–Ascoli Conjecture that could be proven using “addresses” in Section 4.3.
Acknowledgements: We would like to thank Anna Sakovich (Uppsala) for online conversations in July 2025 that lead us to pursuing this project. We all gratefully acknowledge support from the Simons Center for Geometry and Physics and Stony Brook University where we had the opportunity to meet together in person in September 2025 as part of the Program Geometry and Convergence in Mathematical General Relativity. R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. C. Sormani was funded in part by a PSC-CUNY grant. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ 2.2. Review of Gromov–Hausdorff Distance
Gromov defined his intrinsic Hausdorff distance between metric spaces, $(X_{j},d_{j})$, which do not lie in a common metric space in [18]. This is now called the Gromov–Hausdorff (GH) distance. The GH distance is intrinsic in the sense that it depends only on the intrinsic geometry of the two given metric spaces and not the extrinsic geometry of some common ambient space.
In order to define this distance intrinsically, Gromov considered all possible common metric spaces, $(Z,d_{Z})$, and considered maps: $\varphi_{j}\colon X_{j}\to Z$ that are distance preserving
(7) $d_{Z}(\varphi_{j}(p),\varphi_{j}(q))=d_{j}(p,q)\qquad\forall p,q\in X_{j}.$
and takes the infimum:
@@
Thanks to the infimum taken over all possible distance preserving maps, the Gromov–Hausdorff distance depends only on the intrinsic geometry of the original pair of metric spaces not on any particular external metric space, $(Z,d_{Z})$. Yet at the same time the GH distance between metric spaces has many of the same properties as the Hausdorff distance. In particular, there are correspondences between $X_{1}$ and $X_{2}$ with controlled distance distortion that are often used to define the GH distance, but we will not use correspondences in this paper.
@@ 2.4. Review of Fréchet Maps
Given a separable metric space, $(X,d)$, for any countable dense collection of points $\mathcal{N}=\{x_{1},x_{2},$...\} of $X$, one can define a Fréchet map
@@ 2.6. Review of Timed-Fréchet Maps
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau\colon X\to[0,\tau_{max}]$ and any choice of countably dense points in $X$.
@@
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau\colon X\to[0,\tau_{max}]$ that encodes causality as in (23).
@@ 3. Proving the Compactness Theorem with Addresses
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed metric spaces that are equibounded,
@@
and have uniform bounds on their Lipschitz one functions
@@
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, up to a subsequence, there exist distance and time preserving timed-Fréchet maps,
@@
In addition, there is an uncountable index set $\mathcal{A}$, and surjective index maps,
@@ 3.1. Selection of Balls in a Compact Metric Space
Suppose $X_{j}$ is a metric space satisfying the equiboundedness and equicompactness conditions as in the statement of Theorem 2.1. Take the sequence $\varepsilon_{i}=2^{-i}$ and let $N_{i}=N(\varepsilon_{i})$ be natural numbers such that $X_{j}$ can be covered by $N_{i}$ balls of radius $\varepsilon_{i}$. Then we can find a countable dense subset
(47) $\{x_{a}^{j}:\,a\in A\}\subset X_{j}$
using a uniform index set $A$, not depending on $j$, as described below:
For each $i\in{\mathbb{N}}$, we let $A_{i}$ be the finite set,
(48) $A_{i}=\Bigl\{(a_{1},a_{2},\ldots,a_{i}):\,1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\dots,1\leq a_{i}\leq N_{i}\Bigr\},$
and we define the projection maps,
(49) $p_{i}\colon A_{i+1}\to A_{i}\qquad p_{i}\colon(a_{1},\ldots,a_{i+1})\mapsto(a_{1},\ldots,a_{i}).$
Then we define the countable index set for our countable collection of points as
@@
We will define maps
@@
where $I^{j}$ is given by
(52) $I^{j}(a)=I^{j}_{i}(a)\in X_{j}\qquad\forall a\in A_{i}\subset A,$
and set
(53) $x^{j}_{a}=I^{j}(a)=I^{j}_{i}(a)\qquad\forall a\in A_{i},$
which will be well chosen centers of balls of radius $\varepsilon_{i}$ in $X_{j}$ for $a\in A_{i}$. Specifically, first $I^{j}_{i}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, which means
(54) $X_{j}\subset\bigcup_{a\in A_{i}}B_{d_{j}}(x^{j}_{a},\varepsilon_{i})$
Second, for each $a\in A_{i+1}$,
(55) $I^{j}_{i+1}(a)\in B_{d_{j}}(I^{j}_{i}(p_{i}(a)),2\varepsilon_{i})\subset X_{j}$
@@
Hence, the collection of points
(57) ${\mathcal{N}}_{j}=I^{j}(A)=\{x_{a}^{j}:\,a\in A\}=\bigcup_{i\in{\mathbb{N}}}I^{j}_{i}(A_{i})\subset X_{j}$
is countably dense in $X_{j}$ and has been carefully indexed to satisfy the previous two properties.
@@ 3.2. The Index Set of Addresses
A natural consequence of our description of Gromov’s Proposition 3.2, is that we have a uniform way of providing an “addresses” to all points in equicompact collections of metric spaces. More precisely, we have the following definition and proposition.
@@
Let
(80) $q_{i}\colon{\mathcal{A}}\to A_{i}\textrm{ s.t. }q_{i}(\alpha)=\alpha_{i}.$
Then for any $X_{j}$ and any maps $I^{j}_{i}\colon A_{i}\to X_{j}$ satisfying (51)- (55) of Proposition 3.2, the map, ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$, defined by
@@ 3.3. Compactness Theorem with Addresses
Since the time functions are Lipschitz one, we have
@@
which is well defined by (104) and is Lipschitz one by (102).
@@
and a bounded time function, $\tau_{\infty}$, extended as a Lipschitz one function from $X_{\infty}$ to $\bar{X}_{\infty}$.
@@
Note that we have proven that $\bar{X}_{\infty}$ satisfies the conclusions (51)-(57), of Proposition 3.2. It satisfies 54 by (111). Moreover, it easily satisfies (55) as follows. Let $a\in A_{i+1}$,
@@
Let $\epsilon=\varepsilon/3$ and $i\in\mathbb{N}$ such that $\varepsilon_{i}<\epsilon$. Then there exists $N_{\epsilon}^{i}\in{\mathbb{N}}$, as in (125), such that for all $j\geq N_{\epsilon}^{i}$ we have (126).
@@
(140) $\leq\,\,$ $\,\,\varepsilon_{i}+\epsilon+\varepsilon_{i}\,\,\leq\,\,\varepsilon\qquad\forall j\geq N^{i}_{\epsilon},\quad\quad\quad$
where for the second and third term in the right hand side we used that $\tau_{j},\tau_{\infty}$ are Lipschitz one and (83), and for the term in the middle we use (126). Similarly, we get the second inequality,
@@
(147) $\,\,\leq\,\,$ $\,\,\varepsilon_{i}+\epsilon+\varepsilon_{i}\leq\varepsilon\qquad\forall j\geq N^{i}_{\epsilon},\quad\quad\quad$
where we used that $\kappa_{j},\kappa_{\infty}$ are Lipschitz one, (83) and (126).
@@ 4.1. Gromov’s Compactness Theorem with Addresses
If $(X_{j},d_{j})$ is a sequence of compact metric spaces that are equibounded as in (9) and equicompact as in (10) then a subsequence converges in the Gromov–Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty})$. In fact, there exist distance preserving Fréchet maps,
@@
In addition, there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
@@
Take $\tau_{j}$ to be constant zero valued function. Then we satisfy the hypotheses of Theorem 3.1. So the conclusions hold. Finally apply

### v3

@@ Abstract.
The intrinsic timed-Hausdorff distance between timed-metric-spaces, first introduced by Sakovich–Sormani, yields a weak notion of convergence for space-times. In this paper we prove a compactness theorem for the intrinsic timed-Hausdorff convergence of timed-metric-spaces using timed-Fréchet maps. Our proof introduces the notion of “addresses” and provides a new way of stating Gromov’s original compactness theorem for Gromov–Hausdorff convergence of metric spaces. We also obtain a new Arzelà–Ascoli theorem for real valued uniformly bounded Lipschitz functions on Gromov–Hausdorff converging compact metric spaces. Moreover, we establish the triangle inequality for the intrinsic timed-Hausdorff distance.
@@ 1. Introduction
One would also like to study sequences of smooth Lorentzian manifolds which are not diffeomorphic. In [26], Sakovich and Sormani introduced the notion of the intrinsic timed-Hausdorff convergence as one possible way to achieve this. They described a method of converting smooth causally null compactifiable space-times into compact timed-metric-spaces in a canonical way using the cosmological time function of Anderson–Galloway–Howard [4] and the null distance of Sormani–Vega [30]. In general, timed-metric-spaces, $(X_{j},d_{j},\tau_{j})$, are metric spaces, $(X_{j},d_{j})$, endowed with a $1$-Lipschitz function, $\tau_{j}\colon X_{j}\to\mathbb{R}$, and a causal structure. They then defined the intrinsic timed-Hausdorff convergence of these timed-metric-spaces.
@@
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed-metric-spaces that are equibounded,
@@
and have a uniform bound on their $1$-Lipschitz functions
@@
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed-metric-space $(X_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance and time preserving timed-Fréchet maps,
@@
We remark that there are special classes of space-times where the cosmological time functions are easy to control. These include the big-bang space-times and future developed space-times. Sormani–Vega [29] and Sakovich–Sormani [26] defined convergence for such space-times based on the notions of pointed GH convergence [18] and GH convergence of metric pairs by Che, Galaz-García, Guijarro, and Membrillo Solis [7]. Such notions already have compactness theorems proven by Gromov [18] and by Che and Ahumada Gómez [1] respectively. The intrinsic timed-Hausdorff distance applies to a far wider class of space-times including sub-space-times that exhaust asymptotically flat space-times [26].
It should be noted that some authors have used the Sormani–Vega null distance of [30] to convert space-times into metric spaces without keeping track of the time functions. See the work of Allen, Burtscher, García-Heveling, Kunzinger, and Steinbauer in [2, 3, 6, 20]. Furthermore, Allen, Burtscher, Kunzinger, and Steinbauer [2, 3, 20] have applied Gromov’s original compactness theorem to obtain a GH limit space which is only a metric space with no causal structure unlike our timed-metric-spaces.
@@
Our notion of “addresses” provides a new way of stating Gromov’s Compactness Theorem for sequences of metric spaces, that we state and prove in Section 4.1. We also restate an Arzelà–Ascoli Theorem for sequences of real valued uniformly bounded Lipschitz functions on a converging sequence of metric spaces using “addresses” in Section 4.2. We suggest a more general Arzelà–Ascoli Conjecture that could be proven using “addresses” in Section 4.3. We finalize the paper establishing in the Appendix Appendix A: Triangle Inequality for the Intrinsic Timed-Hausdorff Distance the triangle inequality for the intrinsic timed Hausdorff distance.
Acknowledgements: We would like to thank Anna Sakovich (Uppsala University) for online conversations in July 2025 that lead us to pursuing this project. We all gratefully acknowledge support from the Simons Center for Geometry and Physics and Stony Brook University where we had the opportunity to meet together in person in September 2025 as part of the Program Geometry and Convergence in Mathematical General Relativity. We would also like to thank Davide Carazzato (University of Vienna) for sharing ideas on the proof of the triangle inequality for the intrinsic timed-Hausdorff distance. R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. C. Sormani was funded in part by a PSC-CUNY grant. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ 2.2. Review of Gromov–Hausdorff Distance
Gromov defined an intrinsic Hausdorff distance between metric spaces, $(X_{j},d_{j})$, which do not lie in a common metric space in [18]. This is now called the Gromov–Hausdorff (GH) distance. The GH distance is intrinsic in the sense that it depends only on the intrinsic geometry of the two given metric spaces and not the extrinsic geometry of some common ambient space.
In order to define the Gromov-Hausdorff distance between two compact metric spaces, $(X_{j},d_{j})$, consider all possible common metric spaces, $(Z,d_{Z})$, and all maps: $\varphi_{j}\colon X_{j}\to Z$ that are distance preserving,
(7) $d_{Z}(\varphi_{j}(p),\varphi_{j}(q))=d_{j}(p,q)\qquad\forall p,q\in X_{j},$
and take the following infimum:
@@
Thanks to the infimum taken over all possible distance preserving maps, the Gromov–Hausdorff distance depends only on the intrinsic geometry of the original pair of metric spaces. Yet at the same time the GH distance between metric spaces has many of the same properties as the Hausdorff distance. In particular, there are correspondences between $X_{1}$ and $X_{2}$ with controlled distance distortion that are often used to define the GH distance, but we will not use correspondences in this paper.
@@ 2.4. Review of Fréchet Maps
Given a separable and bounded metric space, $(X,d)$, for any countable dense collection of points $\mathcal{N}=\{x_{1},x_{2},$...\} of $X$, one can define a Fréchet map
@@
We remark that one can define Fréchet maps for unbounded metric spaces by fixing a reference point $x_{0}\in X$ and setting
(18) $\kappa_{X}(x)=(d(x_{1},x)-d(x_{1},x_{0}),d(x_{2},x)-d(x_{2},x_{0}),\ldots)\in\ell^{\infty}.$
However, since our main result, Theorem 1.1, concerns compact timed-metric-spaces, Definition 2.2 suffices for our purposes.
@@ 2.6. Review of Timed-Fréchet Maps
for any separable, bounded metric space, $(X,d)$, with a $1$-Lipschitz time function, $\tau\colon X\to[0,\tau_{max}]$ and any choice of countably dense points in $X$.
@@
for any separable, bounded metric space, $(X,d)$, with a $1$-Lipschitz time function, $\tau\colon X\to[0,\tau_{max}]$ that encodes causality as in (24).
@@ 3. Proving the Compactness Theorem with Addresses
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed-metric-spaces that are equibounded,
@@
and have a uniform bounds on their $1$-Lipschitz functions
@@
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed-metric-space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, up to a subsequence, there exist distance and time preserving timed-Fréchet maps,
@@
In addition, there is an uncountable set $\mathcal{A}$, and surjective maps,
@@ 3.1. Selection of Balls in a Compact Metric Space
Suppose $(X_{j},d_{j})_{j}$ is a sequence of metric spaces satisfying the equiboundedness and equicompactness conditions as in the statement of Theorem 2.1. Take the sequence $\varepsilon_{i}=2^{-i}$ and let $N_{i}=N(\varepsilon_{i})$ be natural numbers such that $X_{j}$ can be covered by $N_{i}$ balls of radius $\varepsilon_{i}$. For each $i\in{\mathbb{N}}$, let $A_{i}$ be the finite set,
(48) $A_{i}=\Bigl\{(a_{1},a_{2},\ldots,a_{i}):\,1\leq a_{j}\leq N_{j},j=1,\ldots,i\Bigr\},$
$p_{i}:A_{i+1}\to A_{i}$ the projection maps $p_{i}(a_{1},\ldots,a_{i+1})=(a_{1},\ldots,a_{i})$ and
@@
Then there exists a countable dense subset
(50) ${\mathcal{N}}_{j}=\{x_{a}^{j}:\,a\in A\}\subset X_{j}$
and maps
@@
given by
(52) $I^{j}(a)=I^{j}_{i}(a)=x_{a}^{j}\in X_{j}\qquad\forall a\in A_{i}\subset A,$
that satisfy the following conditions:
Firstly, each $I^{j}_{i}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, which means
(53) $X_{j}\subset\bigcup_{a\in A_{i}}B_{d_{j}}(x^{j}_{a},\varepsilon_{i}).$
Secondly, for each $i,j\in\mathbb{N}$ and $a\in A_{i+1}$,
(54) $I^{j}_{i+1}(a)\in B_{d_{j}}(I^{j}_{i}(p_{i}(a)),2\varepsilon_{i})\subset X_{j},$
@@ 3.2. The Index Set of Addresses
A natural consequence of Proposition 3.2, is that we have a uniform way of providing an “addresses” to all points in equicompact collections of metric spaces. More precisely, we have the following definition and proposition.
@@
Let $q_{i}\colon{\mathcal{A}}\to A_{i}$ given by $q_{i}(\alpha)=\alpha_{i}$. Then for any $X_{j}$ and maps $I^{j}_{i}\colon A_{i}\to X_{j}$ as in Proposition 3.2, the map ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$, defined by
@@ 3.3. Compactness Theorem with Addresses
Since the time functions are $1$-Lipschitz, we have
@@
which is well defined by (101) and is 11-Lipschitz by (99).
@@
and a bounded time function, $\tau_{\infty}$, extended as a $1$-Lipschitz function from $X_{\infty}$ to $\bar{X}_{\infty}$.
@@
Note that we have proven that $\bar{X}_{\infty}$ satisfies the conclusions (51)-(50), of Proposition 3.2. It satisfies 53 by (108). Moreover, it easily satisfies (54) as follows. Let $a\in A_{i+1}$,
@@
Let $\delta=\varepsilon/3$ and $i\in\mathbb{N}$ such that $\varepsilon_{i}<\delta$. Then there exists $N_{\delta}^{i}\in{\mathbb{N}}$, as in (122), such that for all $j\geq N_{\delta}^{i}$ we have (123).
@@
(137) $\leq\,\,$ $\,\,\varepsilon_{i}+\delta+\varepsilon_{i}\,\,\leq\,\,\varepsilon\qquad\forall j\geq N^{i}_{\delta},\quad\quad\quad$
where for the second and third term in the right hand side we used that $\tau_{j},\tau_{\infty}$ are $1$-Lipschitz and (80), and for the term in the middle we use (123). Similarly, we get the second inequality,
@@
(144) $\,\,\leq\,\,$ $\,\,\varepsilon_{i}+\delta+\varepsilon_{i}\leq\varepsilon\qquad\forall j\geq N^{i}_{\delta},\quad\quad\quad$
where we used that $\kappa_{j},\kappa_{\infty}$ are $1$-Lipschitz, (80) and (123).
@@ 4.1. Gromov’s Compactness Theorem with Addresses
If $(X_{j},d_{j})$ is a sequence of compact metric spaces that are equibounded as in (9) and equicompact as in (10) then a subsequence converges in the Gromov–Hausdorff sense to a compact metric space $(X_{\infty},d_{\infty})$. In fact, there exist distance preserving Fréchet maps,
@@
In addition, there is an uncountable set of addresses, $\mathcal{A}$, and surjective index maps,
@@
Take $\tau_{j}$ to be the constant zero valued function. Then all the hypotheses of Theorem 3.1 are satisfied. Thus the conclusions hold. Finally apply